bewellplus.gsu.edu
avataytay03
avataytay03 avataytay03
  • 27-04-2015
  • Mathematics
Answered



7) Jan has two spinners. One has 3 equal sections labeled A, B, and C. The other has 4 equal sections labeled 1, 2, 3, and 4. She spins both spinners. What is the probability that she gets A3?

Verified Answer :

JaySL JaySL
  • 27-04-2015

One out of three results on the first spinner is A, and one out of four on the second is 3.  Thus, the overall probability is 1/3*1/4=1/12.

Answer Link

Related Questions

What is 2009345 in word form
Which property of addition is shown in the equation below abi00i abi
How many times can 23 can go into 1 12
Which of the following symbols in the tragedy of macbeth act ii scene ii signals that the murder has been accomplished
how does blood maintain homeostasis in the body
A Newton is 1 kgms2 Is this a fundamental unit or a derived unit
How does random fertilization add to the genetic variation ANonsister chromatids exchange gene segments to produce new combinations of maternal and paternal gen
the first term in a sequence is an odd number the rule is to multiply by 2 explain why the rest of the terms in the sequence will be even numbers
What is the answer for number six
What tose by mean in mathematics